# High-fidelity universal gates in the 171Yb ground state nuclear spin qubit

### v1

@@ Appendix G Gate Simulator and Error Budget
Based on atomic experiments, we can estimate the contribution of known error sources. We use an approach where we compute our CZ gate using a super-operator simulator with measured experimental values as inputs. Errors are averaged over time, and represent our understanding of the typical operational state of the machine.
@@
We summarize the contributions of each error source with $1-F>10^{-4}$ in Figure 9. Here, the height of each bar represents the decrease in average gate fidelity due to this error. The darker portion of the bar gives the probability for either atom to be outside of the qubit subspace following the CZ gate, i.e. at least one atom is either leaked or lost. The lighter portion alone is the portion of the error not due to leakage or loss. The bars are colored according to whether they occur on the clock operation (green) or the UV operation (magenta). Along with these error sources, we show the experimentally measured CZ CRB and GERB infidelities, as well as the numerically modeled totals for both leakage and loss, and the decrease in fidelity $\Delta F$. This error budget predicts an infidelity of $0.375\%$ with at least one atom leaked or lost $0.211\%$ of the time.

### v2

@@ Appendix G Gate Simulator and Error Budget
Based on atomic experiments, we can estimate the contribution of known error sources. We use an approach where we compute our CZ gate using a super-operator simulator with measured experimental values as inputs. Errors are averaged over time, and represent our understanding of the typical operational state of the machine. This estimation is based on a single gate, rather than a calculation of the fidelity over some specific characterization circuit, i.e. GERB or CRB.
@@
Error source Infidelity Leakage/Loss
Clock detuning 0.126% 0.007%
Clock frequency 0.116% 0.088%
Clock temperature 0.013% 0.013%
Clock scattering 0.019% 0.008%
Clock loss 0.005% 0.005%
Clock light shift 0.009% 0.009%
UV $T_{1}$ 0.075% 0.075%
UV $T_{2}^{*}$ 0.007% 0.001%
UV repeatability 0.007% 0.005%
Totals 0.375% 0.211%
Table 2: Summary of simulated error sources in our CZ gate, as compared to fig. 9.
We summarize the contributions of each error source with $1-F>10^{-4}$ in Figure  9 and also in Table 2. In the figure, the height of each bar represents the decrease in average gate fidelity due to this error. The darker portion of the bar gives the probability for either atom to be outside of the qubit subspace following the CZ gate, i.e. at least one atom is either leaked or lost. The lighter portion alone is the portion of the error not due to leakage or loss. The bars are colored according to whether they occur on the clock operation (green) or the UV operation (magenta). Along with these error sources, we show the experimentally measured CZ CRB and GERB infidelities, as well as the numerically modeled totals for both leakage and loss, and the decrease in fidelity $\Delta F$. This error budget predicts an infidelity of $0.375\%$ with at least one atom leaked or lost $0.211\%$ of the time.